# oMEGACat II - Photometry and proper motions for 1.4 million stars in Omega Centauri and its rotation in the plane of the sky

### v1

@@ 6.3 Creation of a model for CTE effects
As described in Section 4.1, we use CTE corrected input images of the *flc.fits type. However, the applied CTE correction underpredicts the evolution of the CTE loss for the most recent WFC3/UVIS observations. This leads to noticeable CTE effects in both the astrometry and photometry of observations taken after 2017, especially for images with a low background (i.e. those with a short exposure time and those with a very blue filter). Since the start of this work, hst1pass (Anderson, 2022) has been updated with a new and improved CTE correction. However, this correction is meant to be applied to astro-photometry measured on uncorrected *flt.fits type images. Since this improved CTE-correction routine is not currently included in KS2, to our KS2-based single-exposure catalogs we applied empirical corrections based on the comparison between the hst1pass runs on *flt.fits and *flc.fits images. The largest corrections are applied to the UV filters (F275W, F336W) of the GO-16777 program. For the faintest stars ($m_{\rm inst.}\sim-6$) at the largest distances from the amplifier the corrections can reach up to 1 mag, for brighter stars they are much lower.
@@ 6.5.2 Results
We show the statistical properties of the correction in Table 3 and detailed maps and histograms in the Appendix (Figure 16 and 17). The maps show various patterns, that can partially be attributed to physical differential reddening, but also transitions between different pointings. A detailed decomposition into those two components is out of the scope of this work and would not further improve the corrected photometry.
@@ 6.6 Treatment of bright stars
The individual photometric measurements discussed in the section above were all performed with the software KS2. However, this software is not able to measure saturated stars. This limits our completness at the bright end, as the brightest red giant stars are saturated even in the shortest exposures of e.g. the F625W and F814W filters. Including precise photometry for these stars is still important, as they have the highest S/N in spectroscopic studies. We substituted the missing KS2 measurements with data from our single-image hst1pass catalogs, using the same zero-points as determined in the iterative procedure described in Section 6.4.
7 Catalog validation
7.1 Search for systematic effects in magnitude and color
In addition to the local trends in magnitude, we also searched for global trends in magnitude and color, by dividing the proper motions in $m_{\rm F625W}$-magnitude, or $m_{\rm F625W}-m_{\rm F814W}$ color bins (see Figure 18). We calculated the median of the two proper-motion components in each bin and did not see any significant deviation from zero, neither for the raw nor for the a-posteriori corrected proper motions (see Figure 18 in the Appendix).
7.2 Comparison with literature catalogs
7.2.1 Comparison of general properties and completeness
@@
In Figure 12 we compare various general properties of the three different datasets: the spatial coverage, the magnitude-dependent proper-motion errors, the source density as a function of radius and the distribution of magnitudes. We can summarize our findings as follows:
The Bellini et al. (2017a) catalog is limited to a region with only half of the radius of the coverage of our catalog. In this inner region the source density of the photometric catalog is only slightly lower than in our new catalog which is expected, as the Bellini et al. (2017a) catalog is based on a similar dataset and the source detection was performed with the same software. However, if we restrict the comparison to stars where the Bellini et al. (2017a) catalog has proper motion measurements, the density in the literature drops by a factor of around two. This is also expected, as the astrometric part of the Bellini et al. (2017a) catalog is actually based on a previous data reduction (Bellini et al., 2014) with less sensitive photometry software and fewer available data. The proper motion errors of the two HST catalogs show a similar dependence on the magnitude, however, the errors in our new catalog are typically lower by a factor of $\sim$2, due to the significantly longer temporal baseline of our catalog. In addition our proper motion catalog reaches almost 2 magnitudes deeper.
@@
7.2.2 Crossmatch and direct comparison with Bellini et al. (2017a)
We transformed the pixel-based coordinate system of Bellini et al. (2017a) to our own reference system using 6 parameter linear transformations. After that, we used a matching radius of 1 WFC/UVIS pixel (40 mas) to crossmatch stars. This radius is large enough, as the reference epochs of the two catalogs are similar (Bellini et al.: 2007.0; This work: 2012.0). In 5 years the stars will have an RMS displacement of just 0.08 pixel. As it can be seen in Figure 19 in the Appendix, the Bellini et al. catalog is fully contained within our larger field and most stars can be crossmatched (465 362/478 477 for the photometric and 242934/245443 for the astrometric catalog). While the Bellini et al. catalog contains photometry for 26 filters, the focus of our study was on the 6 (3 ACS/WFC, 3 WFC3/UVIS) filters for which we have full coverage out to the half-light radius and the WFC3/UVIS F606W filter for which we have the largest number of photometric measurements in the center. Therefore, there are only 4 filters that we can compare. We show the difference of the method 1 (see Section 4.6) photometry for these 4 filters in Figure 19. The overall agreement is good and only very small systematic shifts of the zero point can be observed (F275W: -0.007; F336W: 0.014; F606W: 0.027; F814W: 0.006). We attribute those differences to the updated instrumental zero point values (see Section 6.1) and the slightly different radii used to create the aperture-photometry-based reference systems.
To directly compare the proper motions, we restrict our analysis to stars brighter than $m_{\rm F606W}=18$, as their statistical errors are lower and potential systematic effects are easier to detect. Figure 20 in the appendix shows the comparison of the raw proper motions. While the overall agreement is good (RMS of difference $\sim$0.09 mas yr-1 in both components), one can see some low spatial frequency effects with amplitudes of up to $\sim$0.1 mas yr-1. These systematic deviations cannot be explained using the individual proper-motion errors alone (the error distributions are 2.31/2.56 times wider than what would be expected from the proper motions errors alone). This is not unexpected for the raw proper motions and can be attributed to CTE effects and residual distortion and is also why we employed the local a-posteriori corrections (see Section 5.3).
In Figure 21 in the appendix, we compare the proper motions after the local a-posteriori corrections have been applied in both catalogs. As expected, the low spatial frequency pattern has disappeared. Instead, we now can see some granularity which is most likely caused by the spatial scale of the local corrections and the limited number of available reference stars. The errors of the corrected proper motions do account for this additional uncertainty and, therefore, the distribution of the residuals is now much more compatible with the errors in the proper-motion catalogs (1.36 / 1.33 times wider than what would be expected from the proper motions errors). The magnitude dependence of the deviations between the proper-motions from the two catalogs is compared in Figure 22 in the appendix, as expected the deviations increase for fainter stars.
7.2.3 Crossmatch and direct comparison with Gaia DR3 & FPR
As already described above, the Gaia DR3 is affected by crowding and has been extended by the Gaia FPR. Here, we crossmatch the combined DR3 + FPR Gaia catalog with our new HST catalog. Using a simple geometric cutoff of 160 mas, we find 353 403 stars in common with our catalog (26 515 from DR3 and 326 888 from the FPR). As many of the crossmatched Gaia sources have limited accuracy and are affected by crowding, we limited ourselves to a subset of well-measured stars. We used a simple proper motion error cut of 0.6 mas yr-1 in both components and both datasets. In addition, we restrict the matching radius to 40 mas (1 WFC3/UVIS pixel) for the high-quality subset. This high-quality (HQ) subset of crossmatched stars contains 7 558 entries from DR3 and 24 447 from the FPR. In Figure 13 (left), it can be seen that, while the actual stellar density increases towards the center, the number of well-measured Gaia stars decreases. Also, it can be seen in the color-magnitude diagram in Figure 13 (middle panel) that the overlap between the catalogs is limited to a relatively small magnitude range of mostly evolved stars. In the right panel of Figure 13, we study the positional residuals between our catalog and well-measured stars from the Gaia catalogs. Given the different reference epochs of the catalogs (our catalog: 2012.0, Gaia DR3: 2016.0; Gaia FPR: 2017.5), and the fact that we anchored our (co-moving) astrometric reference system on observations from 2002.5, we expect both a systematic shift (based on the absolute cluster motion with respect to 2002.5), and random deviations (caused by the individual stars random motion). As it can be seen in the figure, the absolute deviation is around 101 mas for DR3, and 112 mas for FPR, which is compatible with what we expect from the known absolute cluster motion from our initial 2002.5 epoch towards the respective Gaia reference epochs. The random position deviation has an RMS of $\sim$ 3 mas (DR3) and 4 mas (FPR), also compatible with the displacement expected from the velocity dispersion of the cluster.
Figure 23 (DR3) and 24 (FPR) in the appendix show the direct comparison and the differences between the proper motion components between the Gaia datasets and our new catalog. While there is an overall good agreement in terms of pure proper-motion values after the absolute proper motion of $\omega$ Cen is accounted for, there is one large-scale systematic trend that leads to differences of up to $\sim$0.3 mas yr-1. The reason for these systematic differences lies in the different approaches used to measure proper motions: While Gaia measures absolute proper motions anchored to a fixed reference frame, we measure our HST proper motions relative to the bulk motion of cluster stars. We further discuss these systematic differences between the two catalogs in the following Section 8, where we use them to obtain a new, accurate measurement of $\omega$ Cen’s rotation curve.
@@ 8.1 Measuring the rotation curve
By combining our new relative proper-motion catalog with the absolute measurements from Gaia DR3 & FPR, we can overcome the limitations of these past works: By calculating the difference between the absolute Gaia proper motion and our relative proper motions, we obtain a direct measurement of the bulk-motion and any local proper motion trends. This works, as the Gaia proper motions can be seen as a superposition of the bulk motion of the cluster, locally varying systematic motion such as rotation, and the random motion of individual stars, while our relative proper motions only contain the random motions relative to the bulk motion. We study these differences both in 2 dimensional maps and radially. To this aim, we used the Voronoi binning code of Cappellari & Copin (2003) to create two-dimensional bins containing $\sim$250 sources each, and manually created radial bins of $0.5\arcmin$ width. The results, displayed in Figure 14, show a clear rotation pattern with a gradual increase of the rotation in the inner 2 arcminutes flattening out at an amplitude of $\sim$0.3 mas yr-1. At an assumed cluster distance of 5.43 kpc (Baumgardt & Vasiliev, 2021) this corresponds to a rotation of around 6 km s-1, similar to what has been observed using MUSE line-of-sight velocity data (Kamann et al., 2018). Other proper motion studies measure a similar rotation amplitude, but were typically limited to regions at larger distances from the center (van Leeuwen et al., 2000; van Leeuwen & Le Poole, 2002; van de Ven et al., 2006; Bianchini et al., 2018; Vasiliev & Baumgardt, 2021).
Figure 14: Plane of sky rotation determined in 2-dimensional Voronoi bins (left) and as radial profiles (right). The rotation profile from Vasiliev & Baumgardt (2021) shows a significantly shallower increase in the rotation curve. However, it is based on an interpolation of the rotation data towards the center and there were no measurements available to constrain the rotation at smaller radii.
@@ 8.2 Measurements of $\omega$ Cen’s inclination
The linear fit included in the right panel of Fig. 15 corresponds to $D\tan i=5.2\pm 0.3~{}{\rm kpc}$. If we adopt again a distance of $D=5.43~{}{\rm kpc}$, we obtain an inclination of $i=(43.6\pm 1.5)^{\circ}$. This value is in good agreement with previous estimates of the inclination of $\omega$ Cen. van de Ven et al. (2006) derived a value of $i=48\,(+9\,-7)^{\circ}$, while Sollima et al. (2019) found $i=(39.2\pm 4.4)^{\circ}$.
@@ 9 Released data products and recommended use
We make our catalog public in the form of fits tables and Machine Readable ASCII files [^11] A permanent link to download the catalogs will be added once the paper is accepted by the journal. We also aim to submit the tabular tada to the Vizier archive (https://cdsarc.cds.unistra.fr/vizier/submit.htx) . In the following section, we describe the content of the different published files.
@@ 9.1 Astrometric Catalog
We publish a single table that contains the astrometric data including precise positions, proper motions, and several diagnostic parameters. For a description of the different columns see Table E. Our catalogs contains all sources which were recovered in at least two epochs (see Section 5.1). For a small fraction of these sources, no proper motion measurement was possible (either due to saturation or because the proper motion fit did not converge). For these sources we only report the measured position.
@@ 9.1.3 Crossmatches with other catalogs
The column gaia_id contains the gaia source ID of stars from a crossmatch with both Gaia FPR & DR3 (the column gaia_origin specifies the data release). Due to the larger astrometric errors in the Gaia catalogs, we used a matching radius of 120 mas. We did not apply a photometric cut, however this can be used to further refine the selection. Again we only kept the closest cross-match in the case of multiple sources within the matching radius. In total 328 083 stars match these criteria. In addition, we add the flag gaia_hq_subset for the high quality subset that has been used for the rotation curve determination.
@@ 9.2.1 Recommended usage of corrections and errors
We publish a table with the photometric information for each of the 3 ACS/WFC filters (F435W, F625W, F658N) and the 4 WFC3/UVIS filters (F275W, F336W, F606W, F814W). For a description of the different columns see Table E. In general, we recommend the use of our empirical photometric corrections (Section 6.5.1), although this slighlty reduces the coverage.
@@
[…]
@@ 9.2.4 Exemplary photometric correction
Table 4: Number of stars in each filter different that are saturated ($N_{\rm Sat.}$) or are matching our exemplary quality criterion ($N_{\rm HQ}$) comopared to the total number of measurements available in that filter $N_{\rm Total}$
@@ 9.3 Stacked images
Along our astro-photometric catalog, we publish stacked images for the 7 filters, for which we release photometric information. [^12] A permanent link to download these images will be added once the paper is accepted by the journal. The stacked images are normalized to the typical exposure time for the respective filter and combine images from all epochs. Note that the exact flux distribution of sources in the individual images is not preserved in the stacked images, and therefore their main use should be as a high-quality representation of the scene rather than for PSF fitting. The images contain precise WCS information in their header and are also compatible with the pixel coordinates in our astrometric catalog (apart from a shift of $[5000,5000]$ pixels to allow for a smaller image size). An RGB image based on the filters WFC3/UVIS F275W, F336W and F814W can be found in Figure 2.

### v2

@@ 6.3 Creation of a model for CTE effects
As described in Section 4.1, we use CTE corrected input images of the *flc.fits type. However, the applied CTE correction underpredicts the evolution of the CTE loss for the most recent WFC3/UVIS observations. This leads to noticeable CTE effects in both the astrometry and photometry of observations taken after 2017, especially for images with a low background (i.e. those with a short exposure time and those with a very blue filter). Since the start of this work, hst1pass (Anderson, 2022) has been updated with a new and improved CTE correction. However, this correction is meant to be applied to astro-photometry measured on uncorrected *flt.fits type images.
Since this improved CTE-correction routine is not currently included in KS2, to our KS2-based single-exposure catalogs we applied empirical corrections based on the comparison between the hst1pass runs on *flt.fits and *flc.fits images. To derive a model that can transfer the new corrections to our data, we first grouped the data in sets of the same filter and the same exposure time (ensuring the same background level in each such group). After this, we collected the residuals between the hst1pass-flt results and the KS2-flc results. We then modeled these residuals in an instrumental magnitude versus distance-to-amplifier space, to be able to calculate a correction for each measurement in our KS2 based catalogs.
The largest corrections are applied to the UV filters (F275W, F336W) of the GO-16777 program. For the faintest stars ($m_{\rm inst.}\sim-6$) at the largest distances from the amplifier the corrections can reach up to 1 mag, for brighter stars they are much lower.
@@ 6.5.2 Results
We show the statistical properties of the correction in Table 3 and detailed maps and histograms in Appendix B (Figure 16 and 17). The maps show various patterns, that can partially be attributed to physical differential reddening, but also transitions between different pointings. A detailed decomposition into those two components is out of the scope of this work and would not further improve the corrected photometry.
@@ 6.6 Treatment of bright stars
The individual photometric measurements discussed in the section above were all performed with the software KS2. However, this software is not able to measure saturated stars. This limits our completeness at the bright end, as the brightest red giant stars are saturated even in the shortest exposures of e.g. the F625W and F814W filters. Including precise photometry for these stars is still important, as they have the highest S/N in spectroscopic studies. We substituted the missing KS2 measurements with data from our single-image hst1pass catalogs, using the same zero-points as determined in the iterative procedure described in Section 6.4.
7 Comparison with literature catalogs and validation
To validate our new astro-photometric measurements, we performed a search for residual color and magnitude trends (Appendix C) as well as extensive comparisons with previously published HST and Gaia based catalogs. These are described in detail in Appendix D. In the following, we limit ourselves to a comparison of the general catalog properties and a summary of the direct astrometric/photometric comparisons.
7.1 Comparison of general catalog properties and completeness
@@
In Figure 12, we compare various general properties of the three different datasets: the spatial coverage, the magnitude-dependent proper-motion errors, the source density as a function of radius, and the distribution of magnitudes. We can summarize our findings as follows:
The Bellini et al. (2017a) catalog is limited to a region with only half of the radius of the coverage of our catalog. In this inner region the source density of the photometric catalog is only slightly lower than in our new catalog which is expected, as the Bellini et al. (2017a) catalog is based on a similar dataset and the source detection was performed with the same software. However, if we restrict the comparison to stars where the Bellini et al. (2017a) catalog has proper motion measurements, the density in the literature drops by a factor of around two. This is also expected, as the astrometric part of the Bellini et al. (2017a) catalog is actually based on a previous data reduction (Bellini et al., 2014) with less sensitive photometry software and fewer available data. The proper motion errors of the two HST catalogs show a similar dependence on the magnitude, however, the errors in our new catalog are typically lower by a factor of $\sim$2, due to the significantly longer temporal baseline of our catalog. In addition, our proper motion catalog reaches almost 2 magnitudes deeper.
@@
7.2 Crossmatch and direct comparison of measurements
We crossmatched our new catalog with both the Bellini et al. (2017a) and the Gaia literature catalogs to directly compare the measurements. Details of this comparison are described in the Appendix D. To summarize, both catalogs were almost fully included in our new catalogs. With a simple geometric match, we could recover more than 98% of all literature sources, highlighting the completeness of our new catalog, but also the astrometric consistency with the previous works. There is good agreement between both the photometric and the astrometric parts of the Bellini et al. (2017a) catalog.
When comparing our newly measured proper motions with the Gaia catalog, more than 99% of the Gaia sources within our field of view can be recovered in the HST catalog. The differences between the measured positions in both catalogs can be explained with the absolute motion of $\omega$ Cen, but also the internal dispersion (see Figure 13). For the further tests described in Appendix D, we restricted ourselves to a subsample of sources well-measured in both catalogs. This restricts us to relatively bright stars (Figure 13, middle) and also highlights the crowding limitations of the Gaia catalog in the crowded cluster center (Figure 13, left). When comparing the proper motions, one immediately notices the fundamental difference between our locally measured, relative proper motions and the absolute Gaia measurements. The residuals of the comparison can be explained by bulk motion of $\omega$ Cen but also its rotation in the plane of the sky. This is used in the following section to measure the cluster’s rotation.
@@ 8.1 Measuring the rotation curve
By combining our new relative proper-motion catalog with the absolute measurements from Gaia DR3 & FPR, we can overcome the limitations of these past works: By calculating the difference between the absolute Gaia proper motion and our relative proper motions, we obtain a direct measurement of the bulk-motion and any local proper motion trends. This works, as the Gaia proper motions can be seen as a superposition of the bulk motion of the cluster, locally varying systematic motion such as rotation, and the random motion of individual stars, while our relative proper motions only contain the random motions relative to the bulk motion.
We study these differences for a subset of 30 364 well-measured stars (see Appendix D) both in 2-dimensional maps and radially. To this aim, we used the Voronoi binning code of Cappellari & Copin (2003) to create two-dimensional bins containing $\sim$250 sources each, and manually created radial bins of $0.5\arcmin$ width. The results, displayed in Figure 14, show a clear rotation pattern with a gradual increase of the rotation in the inner 2 arcminutes flattening out at an amplitude of $\sim$0.3 mas yr-1. At an assumed cluster distance of 5.43 kpc (Baumgardt & Vasiliev, 2021) this corresponds to a rotation of around 7 km s-1, similar to what has been observed using MUSE line-of-sight velocity data (Kamann et al., 2018). We provide the numerical values of the rotation profile in Appendix F; Table 10 and in a machine readable format. Other proper motion studies measure a similar rotation amplitude, but were typically limited to regions at larger distances from the center (van Leeuwen et al., 2000; van Leeuwen & Le Poole, 2002; van de Ven et al., 2006; Bianchini et al., 2018; Vasiliev & Baumgardt, 2021). This also explains the difference between our measurements and the rotation profile derived in Vasiliev & Baumgardt (2021), see Figure 14. In the innermost region ($r<3\arcmin$), we see a much steeper increase of the rotation profile than described by Vasiliev & Baumgardt (2021), however, this region was previously unconstrained due to the lack of data at small radii.
Figure 14: Plane of sky rotation determined in 2-dimensional Voronoi bins (left) and as radial profiles (right). The numerical values of the rotation profile are also listed in Appendix F; Table 10. In comparison, the rotation profile from Vasiliev & Baumgardt (2021) shows a significantly shallower increase in the rotation curve. However, it is based on an interpolation of the rotation data towards the center and there were no measurements available to constrain the rotation at smaller radii.
@@ 8.2 Measurements of $\omega$ Cen’s inclination
The linear fit included in the right panel of Fig. 15 corresponds to $D\tan i=5.23\pm 0.23~{}{\rm kpc}$. If we adopt again a distance of $D=(5.43\pm 0.05)~{}{\rm kpc}$ (Baumgardt & Vasiliev, 2021), we obtain an inclination of $i=(43.9\pm 1.3)^{\circ}$. This value is in good agreement with previous estimates of the inclination of $\omega$ Cen. van de Ven et al. (2006) derived a value of $i=48^{\circ}\,(+9\,-7)^{\circ}$, while Sollima et al. (2019) found $i=(39.2\pm 4.4)^{\circ}$.
@@ 9 Released data products and recommended use
We make our catalog public in the form of fits tables and Machine Readable ASCII files in a repository hosted by Zenodo ( doi:10.5281/zenodo.11104046). In the following section, we describe the content of the different published files.
@@ 9.1 Astrometric Catalog
We publish a single table that contains the astrometric data including precise positions, proper motions, and several diagnostic parameters. For a description of the different columns see Appendix E; Table E. Our catalogs contains all sources which were recovered in at least two epochs (see Section 5.1). For a small fraction of these sources, no proper motion measurement was possible (either due to saturation or because the proper motion fit did not converge). For these sources we only report the measured position.
@@ 9.1.3 Crossmatches with other catalogs
The column gaia_id contains the gaia source ID of stars from a crossmatch with both Gaia FPR & DR3 (the column gaia_origin specifies the data release). Due to the larger astrometric errors in the Gaia catalogs, we used a matching radius of 160 mas. We did not apply a photometric cut, however this can be used to further refine the selection. Again we only kept the closest cross-match in the case of multiple sources within the matching radius. In total 373 291 stars match these criteria. To facilitate further comparisons between our dataset and the Gaia measurements we also include several key quantities from the crossmatched Gaia catalog in our release data products, including the source positions measured with Gaia, the absolute Gaia proper motion and the Gaia G photometry. The Gaia proper motions can be also used to substitute the missing proper motions for stars too bright for HST measurements. In that case, the different definitions of the proper motions (oMEGACat: relative; Gaia: absolute) have to be taken into account. In addition, we add the flag gaia_hq_subset for the high quality subset that has been used for the rotation curve determination.
@@ 9.2.1 Recommended usage of corrections and errors
We publish a table with the photometric information for each of the 3 ACS/WFC filters (F435W, F625W, F658N) and the 4 WFC3/UVIS filters (F275W, F336W, F606W, F814W). For a description of the different columns see Appendix E; Table E. In general, we recommend the use of our empirical photometric corrections (Section 6.5.1), although this slighlty reduces the coverage.
@@
For most photometric use-cases we recommend to use the weighted mean of method 1 (see Section 4.6) photometry (m1_weighted_mean). The corresponding weighted mean error is saved as column m1_weighted_mean_error. This error may be underestimated in cases of crowding, therefore, we recommend scaling it with the square-root of the chi2_red whenever $\chi^{2}_{red.}>1.0$. When applying the empirical photometric corrections determined in Section 6.5.1 (which we also recommend for most cases), one additionally has to add the error on the correction (Table 3) in quadrature. For the convenience of the user, we provide the corrected photometry and the combined error in the first two columns corrected_mag and corrected_mag_error in the published data products. We also note here that the absolute zeropoints have reported uncertainties of $\sim$1(ACS, Bohlin 2016) and 2-3% (UVIS, Calamida et al. 2022), which corresponds to absolute uncertainties at the 0.02-0.03 mag level. This absolute uncertainty does not affect the internal consistency of our catalog (which is ensured by the corrections), but has to be taken into account when comparing the data e.g. with isochrone models.
@@ 9.2.4 Exemplary photometric correction
Table 4: Number of stars in each filter different that are saturated ($N_{\rm Sat.}$) or are matching our exemplary quality criterion ($N_{\rm HQ}$) compared to the total number of measurements available in that filter $N_{\rm Total}$
@@ 9.3 Stacked images
Along our astro-photometric catalog, we publish stacked images for the 7 filters, for which we release photometric information. The stacked images are normalized to the typical exposure time for the respective filter (see Table 5) and combine images from all epochs. Note that the exact flux distribution of sources in the individual images is not preserved in the stacked images, and therefore their main use should be as a high-quality representation of the scene rather than for PSF fitting. The images contain precise WCS information in their header and are also compatible with the pixel coordinates in our astrometric catalog (apart from a shift of $[5000,5000]$ pixels to allow for a smaller image size). An RGB image based on the filters WFC3/UVIS F275W, F336W and F814W can be found in Figure 2.
Table 5: Exposure times to which the published stacked images are normalized.
Instrument Filter Exposure time
ACS/WFC F435W 340 s
ACS/WFC F625W 340 s
ACS/WFC F658N 440 s
WFC3/UVIS F275W 773 s
WFC3/UVIS F336W 475 s
WFC3/UVIS F606W 40 s
WFC3/UVIS F814W 250 s